# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ 5.1 Experimental Settings
Table 1: Recovery ratios on 6 functionalities under FT-all (SCARA ||||SAP-DP|||| Fully-closed). “H.E.” in Code domain presents the benchmark “HumanEval”. More details are available in Appendix C.2.
Benchmark Llama2-70B Llama2-7B Mistral-7B Phi-2 Phi-1.5
Rsn. PIQA 62.6||||59.8||||63.0 64.7||||64.7||||64.6 63.0||||61.2||||60.2 68.3||||65.6||||65.7 70.6||||69.5||||66.7
Winogrande 68.5||||67.7||||68.3 76.8||||74.8||||76.6 67.2||||69.0||||68.3 68.3||||64.9||||64.8 70.3||||67.8||||67.6
ARC-easy 31.9||||32.8||||31.3 36.3||||35.5||||34.9 32.3||||34.7||||32.0 43.2||||35.3||||33.9 40.5||||37.8||||36.1
ARC-challenge 38.5||||38.1||||44.2 47.8||||46.6||||50.9 39.7||||42.6||||44.5 36.8||||36.6||||35.3 46.1||||44.4||||47.5
Hellaswag 31.4||||31.4||||32.4 33.9||||34.0||||35.0 32.2||||32.0||||31.3 37.4||||37.3||||34.3 42.0||||41.0||||40.0
Read. LAMBADA 0.01||||0.00||||0.00 0.02||||0.00||||0.01 0.16||||0.00||||0.01 1.34||||0.04||||0.00 1.37||||0.00||||0.00
BoolQ 47.2||||47.1||||53.9 59.5||||56.0||||65.0 48.3||||46.8||||56.7 56.7||||50.3||||55.8 61.7||||54.9||||60.8
SQuADv2-EM 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
SQuADv2-F1 1.50||||1.68||||0.34 0.68||||0.88||||0.82 1.69||||0.36||||0.93 3.65||||0.39||||0.90 1.28||||1.07||||2.64
OBQA 54.5||||54.5||||57.1 57.4||||52.5||||59.2 57.7||||56.8||||56.3 0.00||||0.00||||0.02 0.04||||0.00||||0.00
Knl. NaturalQuestions 0.00||||0.02||||0.00 0.01||||0.01||||0.08 0.00||||0.00||||0.02 0.01||||0.00||||0.06 0.21||||0.00||||0.00
TriviaQA 0.00||||0.02||||0.00 0.00||||0.00||||0.03 0.00||||0.00||||0.01 0.01||||0.00||||0.01 0.01||||0.00||||0.00
Code MBPP&H.E. 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
Math GSM8K 0.02||||0.00||||0.06 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00
Gen. MMLU 36.8||||38.3||||36.5 52.9||||50.0||||53.3 40.4||||36.9||||37.2 42.6||||40.3||||40.5 56.7||||54.1||||54.1
BBH 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.00||||0.00||||0.00 0.01||||0.00||||0.00 0.00||||0.00||||0.00
Average Recovery Ratio($\downarrow$) 21.9||||21.8||||22.8 25.3||||24.4||||25.9 22.5||||22.4||||22.8 23.9||||22.3||||22.4 26.2||||25.3||||25.4
Closed-source Ratio($\downarrow$) 2.50||||92.5||||100. 3.16||||81.3||||100. 3.16||||81.6||||100. 6.25||||81.3||||100. 8.33||||75.0||||100.
Table 2: Recovery ratios on Llama2-70B.
Strat. Method Rsn. Read. Knl. C.&M. Gen. ARR
FT-c. SCARA 47.1 21.6 0.00 0.03 18.7 22.6
@@
F-Closed 47.8 21.2 0.00 0.08 18.5 22.8
Table 3: ARR of SCARA vs. attacks.
Model FT-all FT-closed SEM
Llama2-70B 21.9 22.6 22.4
Llama2-7B 25.3 24.8 25.0
Mistral-7B 22.5 22.5 22.0
Phi-2 23.9 23.7 22.1
Phi-1.5 26.2 26.9 24.7
5.2 Main Results
In this subsection, we compare SCARA with three baselines, demonstrating its superior customizability on downstream domains while preserving similar resilience against model recovery attacks.
Customizability: SCARA vs. Baselines. We compare the customization performance of SCARA with closed-source baselines. Results are shown in Figure 4 and detailed in Appendix B.6.
On 70B and 7B models, SCARA consistently surpasses SAP-DP and fully-closed approaches across six domains and aligns closely with the performance of the fully-open approach, where all parameters are accessible. For instance, in the Law domain, SCARA improves scores by 10% over SAP-DP and fully-closed approaches on Llama2-70B, with this improvement rising to 35% on 7B models. Similar patterns of enhanced customizability are also evident in Phi-2 model, though the improvement on the Law domain narrows to only 1%. Furthermore, SCARA maintains performance comparable to the fully-open approach, demonstrating that hiding a small portion of parameters has a small impact on downstream customizability. More results are reported in Appendix C.3.
Resilience: SCARA vs. Baselines. We compare the resilience of SCARA with other closed-source baselines under three recovery strategies: FT-all, FT-closed, and SEM attack. As shown in Table 1, under FT-all attack, SCARA, SAP-DP, and fully-closed approach show similar resilience across various architectures and domains, with ARR differences within 1.4%. For instance, on Llama2-70B, SCARA keeps only 1.25% of parameters hidden, yet achieves an ARR of 21.9%, comparable to SAP-DP (21.8%) and the fully-closed approach (22.8%), which keep 92.5% and 100% of parameters closed-sourced, respectively. This pattern also extends across five architectures, indicating that our approach effectively preserves resilience with limited parameter closed-sourced. As shown in Table 3, SCARA achieves comparable resilience under FT-closed and SEM attacks. Under FT-closed attack, the recovery ratio differences among the three approaches remain under 2.1% across six domains. Similarly, under SEM attack, recovery ratios of SCARA align closely with the other approaches. These results show that SCARA offers resilience against model recovery on par with fully-closed approach, despite requiring fewer private parameters. More details are in Appendix C.4.
Figure 5: Amplification of error in Llama2-7B.
Table 4: SCARA vs. dataset scales.
Scale Rsn. Read. Knl. C.&M. Gen. ARR
51k 51.7 21.6 0.01 0.00 28.3 25.3
100k 51.3 21.5 0.13 0.00 29.6 25.3
200k 51.4 21.7 0.11 0.00 29.7 25.2
300k 51.6 21.7 0.11 0.00 30.5 25.5
500k 51.8 22.0 0.09 0.00 30.8 25.8
Resilience: SCARA vs. Recovery Strategies. Table 3 shows that SCARA effectively defends three recovery attack strategies on all models. We observe that SEM, a typical and effective attack for recovering embedding models, does not show a significant boost in recovery performance. This can be because SEM attackers focus on recovering only the proprietary embedding module, while the semi-open model recovery attackers aim at recovering the full functionality of the entire model, including both proprietary and public modules. The targets of these two attackers are different since even small errors in bottom layers can lead to significant output deviations. To see this, we add small perturbations on parameters in the first layer of Llama2-7B model and evaluate the embedding deviation at the output of each decoder layer. Figure 5 shows that the norm of deviation increases as the layer index increases, indicating that small errors are amplified by subsequent layers, leading to large deviations in the final output. Therefore, SCARA closed-sources the first several layers, effectively leveraging this amplification, making the functionality recovery more difficult and ensuring strong resilience against recovery attack. We report details in Appendix B.7 and C.5.
Resilience: SCARA vs. Recovery Dataset Scales. We further evaluate the resilience of SCARA against FT-all by increasing the recovery dataset scale on Llama2-7B to determine if larger datasets would compromise its effectiveness. More details on the attack dataset are in Appendix 7. Table 4 shows the recovery ratio achieved by SCARA under each attack dataset. We observe that increasing the scale of the attack dataset leads to only a mild increase in recovery ratios, indicating a limited impact on SCARA. For instance, recovering with the 500k samples results in only a 0.5% ARR improvement over 51k samples. This suggests that the resilience provided by SCARA cannot be easily compromised by simply increasing the dataset scale. Details are reported in Appendix C.6.
5.3 Analysis on SCARA
In this subsection, we assess the effectiveness of SCARA on four LLMs, showing the resilience transition in LLMs and the strong correlation between recovery difficulty and ARR.
Figure 6: (a) shows customizability and resilience changes in Llama2-7B and Phi-2 when different closed-sourced layer sets. Each point for customizability represents hidden layers from the start up to the indicated set. (b) presents $\Delta$ARR for different closed-sourced parameter quantities and proportions in Llama2-7B. (c) depicts recovery ratio changes ($\Delta$RR) in specific functions of Llama2-7B with varying closed-sourced parameter ratios.
Transition Layer. We investigate the existence of transition layers in resilience against model recovery and customizability in the math domain. Results are in Figure 6(a) and detailed in Appendix C.7
For resilience, we hide same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B and two layers for Phi-2. We observe clear resilience transitions in both models. For example, in Llama2-7B, the transition occurs at the eighth layer set. For sets preceding the eighth, $\Delta$ARR remains close to zero, indicating that concealing any set up to the eighth layer set yields resilience comparable to the fully-closed approach. However, concealing sets beyond the eighth leads to decreased resilience, as indicated by increasing $\Delta$ARR values. In the smaller Phi-2 model, the transition happens earlier at the first layer set, with hiding later sets resulting in reduced resilience. These patterns suggest resilience transitions occur deeper in larger models. Similar trends of transitions in Mistral-7B and Phi-1.5 can be found in Appendix B.8.
For customizability, we conceal varying numbers of layers from the start and fine-tune the open set. We observe the customizability transition in the 7B model. Specifically in Llama2-7B, concealing only the first layer set improves performance by 43%, but this gain drops to 7% when concealing up to the third layer set and eventually fluctuates around 0%. However, we barely observe transitions in the smaller model, with improvements limited to 3%. This suggests that fewer closed-sourced layers enhance customizability, and larger models possess greater customizability than smaller ones.
Resilience vs. Closed-source Size. We explore how closed-sourced set size impacts the model resilience in Llama2-7B. We start from the smallest self-attention module (k_projection) in a decoder layer and extend to the entire model. Results are shown in Figure 6 and detailed in Appendix C.8.
In the general domain (Figure 6(b)), we observe that resilience emerges when a sufficient number of parameters are closed-sourced. Specifically in Llama2-7B, we observe that when the proportion of closed-sourced increases to 3%, resilience emerges. Furthermore, we observe similar patterns in specific domains, yet resilience transitions vary across capabilities as shown in Figure 6(c). For instance, technical skills such as Math show earlier transitions, with resilience emerging at 1% parameters closed-sourced, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, closed-sourcing a small portion of parameters can provide sufficient resilience against model recovery, meanwhile, technical capabilities tend to be more challenging to recover than other domains. Other models are analyzed in Appendix B.9.
Figure 7: (a) presents the Pearson coefficient between recovery ratio (RR) and recovery difficulty (RD) across four models and six domains. (b) and (c) depict the link between $\Delta$ARR and RD for Llama2-7B and OPT-350M.
5.4 discussions
Effectiveness of RD on large models. We assess the efficacy of the recovery difficulty (RD) in estimating the performance of the recovered model. Specifically, we calculate the Pearson and Spearman correlation coefficients between RD and ARR across different capability groups. As shown in Figure 7(a), we observe a negative correlation between the recovery difficulty and average recovery ratio. For example, in Llama2-7B, the Pearson coefficient is consistently below -0.80, reaching as low as -0.98. We observe similar phenomena in other models with varying architectures and sizes, confirming RD as a reliable predictor of recovered model performance and the effectiveness of SCARA. Further analysis and results of Spearman coefficients can be found in Appendix B.10.
[…]
Figure 8: Sensitivity to $\varepsilon$.
Table 5: Performance of SCARA defending adversarial attacks. $\downarrow$ indicates the smaller the better.
[…]
Gold Std. 58.0 43.9 0.00
SCARA 72.3 85.0 26.5
SAP-DP 72.2 83.9 24.9
[…]
Limited Defense against Adversarial Attack. We compare SCARA and SAP-DP in defending against three black-box adversarial attacks on Llama2-7B. Specifically, we apply the membership inference (Fu et al., 2023) (MIA), attribute inference (Staab et al., 2023) (AIA), and prompt injection (Liu et al., 2023) (PIA) attacks to the semi-open models produced by SAP-DP and SCARA. As shown in Table 5, we observe that SAP-DP outperforms SCARA across all three attacks, but still performs worse than the gold standard. This is because SCARA does not introduce additional output perturbation and thus provide limited defense against black-box adversarial attacks. Details can be found in Appendix B.11.
@@ 6 Conclusion
In this paper, we explored finding minimal closed-sourced sets to enhance LLM customizability while preserving their resilience against semi-open model recovery attacks. We theoretically prove that minor errors in bottom decoder layers prior to a transition layer greatly reduce recovery attack success. We introduced SCARA, which selectively closed-sources a small set of layers, achieving superior customizability and comparable resilience to SAP-DP and fully-closed. We empirically investigated the existence of customization and resilience transitions, showed the impact of closed-source size on model resilience, analyzed the effectiveness of our approach, and finally discussed its limitations.

### v2

@@ 5.1 Experimental Settings
Table 1: Distillation ratios on 6 functionalities under FT-all (SOLID||||SAP-DP||||Fully-secured||||DarkneTZ). “H.E.” in Code domain presents the benchmark “HumanEval”. More results and details are available in Appendix C.2.
Benchmark Llama2-70B Llama2-7B Mistral-7B Phi-2
Rsn. PIQA 62.6||||59.8||||63.0||||99.3 64.7||||64.7||||64.6||||99.1 63.0||||61.2||||60.2||||92.2 68.3||||65.6||||65.7||||99.1
Winogrande 68.5||||67.7||||68.3||||98.3 76.8||||74.8||||76.6||||100. 67.2||||69.0||||68.3||||89.5 68.3||||64.9||||64.8||||99.1
ARC-easy 31.9||||32.8||||31.3||||98.5 36.3||||35.5||||34.9||||97.6 32.3||||34.7||||32.0||||86.6 43.2||||35.3||||33.9||||99.5
ARC-challenge 38.5||||38.1||||44.2||||99.2 47.8||||46.6||||50.9||||100. 39.7||||42.6||||44.5||||81.4 36.8||||36.6||||35.3||||99.5
Hellaswag 31.4||||31.4||||32.4||||98.1 33.9||||34.0||||35.0||||96.6 32.2||||32.0||||31.3||||84.6 37.4||||37.3||||34.3||||96.5
Read. LAMBADA 0.01||||0.00||||0.00||||88.6 0.02||||0.00||||0.01||||92.2 0.16||||0.00||||0.01||||67.9 1.34||||0.04||||0.00||||94.6
BoolQ 47.2||||47.1||||53.9||||100. 59.5||||56.0||||65.0||||99.6 48.3||||46.8||||56.7||||97.3 56.7||||50.3||||55.8||||100.
SQuADv2 1.50||||1.68||||0.34||||55.3 0.68||||0.88||||0.82||||59.5 1.69||||0.36||||0.93||||50.7 3.65||||0.39||||0.90||||62.9
OBQA 54.5||||54.5||||57.1||||99.6 57.4||||52.5||||59.2||||94.8 57.7||||56.8||||56.3||||84.0 0.00||||0.00||||0.02||||94.3
Knl. NaturalQuestions 0.00||||0.02||||0.00||||40.1 0.01||||0.01||||0.08||||53.6 0.00||||0.00||||0.02||||31.8 0.01||||0.00||||0.06||||87.4
TriviaQA 0.00||||0.02||||0.00||||72.3 0.00||||0.00||||0.03||||73.8 0.00||||0.00||||0.01||||38.7 0.01||||0.00||||0.01||||68.9
Code MBPP&H.E. 0.00||||0.00||||0.00||||58.6 0.00||||0.00||||0.00||||90.9 0.00||||0.00||||0.00||||40.2 0.00||||0.00||||0.00||||91.1
Math GSM8K 0.02||||0.00||||0.06||||79.6 0.00||||0.00||||0.00||||78.6 0.00||||0.00||||0.00||||31.1 0.00||||0.00||||0.00||||86.2
Gen. MMLU 36.8||||38.3||||36.5||||96.7 52.9||||50.0||||53.3||||110. 40.4||||36.9||||37.2||||81.7 42.6||||40.3||||40.5||||99.5
BBH 0.00||||0.00||||0.00||||93.3 0.00||||0.00||||0.00||||101. 0.00||||0.00||||0.00||||63.3 0.01||||0.00||||0.00||||94.8
Average Distillation Ratio($\downarrow$) 21.9||||21.8||||22.8||||77.9 25.3||||24.4||||25.9||||86.5 22.5||||22.4||||22.8||||73.7 23.9||||22.3||||22.4||||88.9
Secured Ratio($\downarrow$) 2.50||||92.5||||100.||||1.25 3.16||||81.3||||100.||||3.16 3.16||||81.3||||100.||||3.16 6.25||||81.3||||100.||||3.16
Evaluation Benchmarks We assess adaptability on six downstream tasks: Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.).
Metrics. We measure customization through model’s performance on benchmarks. For security, we calculate the “Average Distillation Ratio” (ADR) by averaging the distillation ratios across benchmarks. A lower ADR indicates higher security offered by the secure set.
5.2 Failure in Defense (RQ1)
We evaluate the security of DarkneTZ using three distillation strategies. Based on the results shown in Tables 1 and 2, we have following observations.
Observation 1: DarkneTZ, which secures only the last decoder layer, fails to protect the model against all three attacks (Table 1). DarkneTZ achieves ADRs generally exceeding 73% across four models, significantly higher than the average 24% ADR obtained by the fully-secured approach that offers maximum security. Specifically, on Llama2-70B and Phi-2, the distillation ratio for the Reasoning (Rsn.) task consistently exceeds 98%, indicating near-complete replication of model functionality. On Llama2-7B, the attack achieves a distillation ratio over 100% on the MMLU and BBH, suggesting the distilled model even outperforms the victim on these tasks. Similarly, Table 2 highlights consistent failure patterns against FT-closed and SEM attacks, with ADRs consistently above 75%. This confirms that securing only the top layers is insufficient to prevent the replication of model functionality.
Figure 4: (a) shows the trade-off between security and customization for Llama2-7B and Phi-2 with different placements of same-sized secured sets. (b) shows the trade-off as the secured set size increases from the first decoder layer. Smaller ADR indicates higher security and higher ACC reflects better customizability.
5.3 Security-Customization Trade-off (RQ2)
We conduct two experiments to analyze the impact of secured layer placement and quantity on the trade-off between security and customization. First, we secure one layer in Llama2-7B and two in Phi-2, varying their placement. Second, we incrementally secure both models by adding protected layers, starting from the smallest module (k_project) of the first decoder layer. These models are evaluated under the FT-all distillation attack and customized for the math domain. The results, as shown in Figure 4, lead to the following observations. More results and details can be found in Appendix B.7 and B.8.
Observation 2: Secured layer placement significantly impacts security, consistent with Theorem 1, but has limited effect on customization performance (Figure 4(a)). For Llama2-7B, security transitions at the fourteenth layer, with ADR consistently near 20% for earlier sets, indicating stronger security than protecting later layers. Meanwhile, customization accuracy remains stable across placements, highlighting the advantage of securing pre-transition layers. In contrast, Phi-2 transitions earlier at the first layer set, where only the first set balances security and customization, with later sets reducing security. These results suggest that securing layers before the transition layer optimizes the security-customization trade-off.
Observation 3: Increasing the number of secured layers enhances security but reduces customization (Figure 4(b)). The ADR of Llama2-7B decreases from 85% to 22% after securing an entire decoder layer, indicating improved security. However, customization accuracy drops from 29% to 21% as the number of secured layers increases from one to five, reflecting reduced customization flexibility. A similar trend is observed in Phi-2, suggesting that while increasing the number of secured layers enhances security (lower ADR), it negatively impacts customization flexibility (lower ACC) in both models.
5.4 Effectiveness of Securing Bottom Layers (RQ3)
We compare the security of SOLID with baseline deployments across three distillation strategies. The results lead to the following observations.
Table 2: Distillation ratios of Llama2-70B under FT-closed (FT-c.) and SEM attacks.
Strat. Method Rsn. Read. Knl. C.&M. Gen. ADR
FT-c. SOLID 47.1 21.6 0.00 0.03 18.7 22.6
@@ 5.4 Effectiveness of Securing Bottom Layers (RQ3)
F-Secured 47.8 21.2 0.00 0.08 18.5 22.8
DarkneTZ 98.8 71.2 54.2 66.3 94.1 77.4
Observation 4: SOLID achieves security against model distillation comparable to fully secured protection while securing significantly fewer parameters (Table 1). SOLID attains a similar security level (ADR) to SAP-DP and the fully secured approach across four architectures and multiple domains, securing at most 6.25% of parameters—compared to at least 80% for the others. For instance, on Llama2-70B, SOLID secures only 1.25% of parameters yet achieves an ADR of 21.9%, comparable to SAP-DP (21.8%) and the fully secured approach (22.8%). Additionally, SOLID maintains this security under FT-closed and SEM attacks, with ADR differences from SAP-DP and the fully secured approach remaining below 2.1% across six domains (Table 2). Details are in Appendix C.4 and C.5.
Figure 5: Customization performance comparison of secured models on six downstream tasks.
Observation 5: SOLID approaches the customization flexibility of full parameter fine-tuning, consistently outperforming baseline deployments (Figure 5). We observe that the performance of SOLID consistently matches full parameter fine-tuning across four architectures, with differences within 4%. Additionally, in the Law domain, SOLID improves performance by 10% over SAP-DP and the fully-secured approach on Llama2-70B, with a 35% improvement on the 7B models. Similar gains are evident on Phi-2, though the improvement in the Law domain reduces to 1%. These results demonstrate that securing a small portion of parameters minimally impacts customization, while providing strong protection against distillation attacks. Further results can be found in Appendix B.6 and C.3.
6 Alternative Views
Acknowledging diverse views and potential curiosities regarding LLM deployment design, we objectively examine several alternate viewpoints.
  • Alternative: Enhanced Regulation. Enhanced government data regulations can partially mitigate ethical privacy concerns and strengthen user confidence in data uploads and fine-tuning APIs, thereby promoting industry development. However, they do not eliminate the risk of data breaches and may struggle to address adaptive threats, rendering them insufficient for organizations with stringent privacy requirements, such as government agencies.
  • Alternative: Hybrid Privacy-Aware Deployment. Federated hybrid deployment enables users to deploy early model layers and process sensitive data locally while hosting later layers in the cloud for enhanced protection (Shen et al., 2023). It employs privacy-preserving techniques (e.g., differential privacy) to sanitize inputs before generating intermediate representations, ensuring secure data transmission. This approach preserves customization performance while improving both data privacy and model security. However, supervised fine-tuning still requires data uploads, and intermediate exchanges remain vulnerable to breaches (Zhou et al., 2022), thereby compromising the overall security assurance.
  • Alternative: Advanced Open-Weight Models. The success of DeepSeek-R1 (Guo et al., 2025) has showcased the power of open-weight models, raising expectations for more robust open-source alternatives. Developing open-weight models with strong security alignment could offer a viable solution to the ongoing dilemma between data privacy and model confidentiality. These models, capable of ensuring both usability and robust misuse detection, could address ethical concerns and serve as an alternative to the privacy-preserving deployment methods. However, their adoption may be constrained by limited profitability.
@@ 7 Conclusion
In this position paper, we advocate for a balanced approach to on-premises LLM deployment, addressing the dilemma between data privacy and model confidentiality. Building on our theorem, we propose SOLID, a simple yet effective semi-open solution that efficiently secures a few bottom layers of LLMs. We believe that our work shifts the view from skepticism to opportunity, paving a viable middle path for secure AI deployment that protects user privacy, safeguards vendor intellectual property, and preserves model customization flexibility. While our position may invite both support and critique, we hope it provides meaningful insights and fosters ongoing discourse. If this paper contributes to constructive engagement within the community, it will have fulfilled its purpose.
Impact Statement
The primary objective of this position paper is to advocate for a balanced approach to on-premises LLM deployment, ensuring both data privacy and model security. We introduce an efficient theoretically inspired semi-open deployment framework that mitigates functional extraction risks while enabling customization, providing a preliminary yet effective solution for privacy-sensitive industries. While our work primarily contributes to academic discourse, it also has broader implications for secure AI deployment in sectors where data confidentiality is critical. Ethically, we highlight the potential to simultaneously prevent data leakage, safeguard model security, and mitigate misuse risks. While we foresee no immediate societal consequences requiring specific emphasis, we acknowledge the importance of ongoing ethical considerations and the potential for future societal impacts as this interdisciplinary field evolves.